OpenStudy (anonymous):

Given the function f(x) = 3x + 1, evaluate f(a + 1).

OpenStudy (solomonzelman):

you mean f(x+1) ?

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ f(x) = 3x + 1\LARGE\color{white}{ \rm │ }}\) and \(\normalsize\color{green}{ f(x+1) = 3(x+1) + 1= ?\LARGE\color{white}{ \rm │ }}\) you tell me

OpenStudy (solomonzelman):

well, f(a+1) would be the exact same thing, BUT with ` a ` instead of ` x ` .

OpenStudy (anonymous):

its the same thing as f(x)

OpenStudy (solomonzelman):

well, you are plugging `a+1` EVERYWHERE for x,

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

my answers to pick is 3a + 1 3a + 2 3a + 4

OpenStudy (solomonzelman):

\(\normalsize\color{black}{ f(a+1) = 3(a+1) + 1}\) \(\normalsize\color{black}{ f(a+1) = 3a+3 + 1}\) \(\normalsize\color{black}{ f(a+1) = ??}\) which option do you pick ?

OpenStudy (anonymous):

you have to distribute

OpenStudy (solomonzelman):

you have to add like terms on the right side.

OpenStudy (anonymous):

okay